QueenMidnight16:

please help! A scale drawing of a living room is shown below. The scale is 1 : 40. A rectangle is shown. The length of the rectangle is labeled 6 inches. The width of the rectangle is labeled 4 inches. Show your work to determine the area of the room in square feet.

MidniteSunny:

What is the rectangle that is shown?

QueenMidnight16:

1 attachment
QueenMidnight16:

thats the image

MidniteSunny:

Ah oki well first I think we gotta figure out the length and width of the actual room, using the ratio 1 : 40 where the scaled down model is 1 and the room is 40.

MidniteSunny:

we can put it like this to try to find out the length: 1:40 = 6:x from here we can cross multiply like this: 1/40=6/x

MidniteSunny:

do yk how to cross multiply ?

QueenMidnight16:

i think

MidniteSunny:

oki ! tell me what u get for 1/40=6/x

QueenMidnight16:

remind me how to cross multiple XD

MidniteSunny:

o oki so when we hav the equation 1/40=6/x, we can put it as its cross product by looking at, say, the 1 and seeing that across and down from the 1 (in 1/40) is x (from 6/x), and then across and down from 6 is 40. If you look at it after being written down, it would make an "x" or a cross

MidniteSunny:

and then we would write it out like 1×x=40×6

MidniteSunny:

i hope that made sense >~< but uh yeah so what do u think it is from there ?

MidniteSunny:

@queenmidnight16

QueenMidnight16:

idkkk

QueenMidnight16:

i dont remember how to multiple 1 by x

MidniteSunny:

o well it would just make it x, cuz 1x can b simplified to just x

QueenMidnight16:

ok so x and 240?

MidniteSunny:

yes ! x=240 ! gj !

QueenMidnight16:

is that the answer?

QueenMidnight16:

x=240?

MidniteSunny:

that's the length of the actual room, we gotta set up another equation to find out the width it's the same as the length, but instead of 6 it's 4

QueenMidnight16:

ok

MidniteSunny:

mhm ! tell me wut chu get !

QueenMidnight16:

soooo its 1xX and 40x4?

MidniteSunny:

mhm

QueenMidnight16:

ok

QueenMidnight16:

x=160?

MidniteSunny:

correcto

MidniteSunny:

nowall that's left ish to find ze area of ze room using the length and width u figured out

MidniteSunny:

now all*

QueenMidnight16:

how

MidniteSunny:

Area = length x width

QueenMidnight16:

160x x 240x?

MidniteSunny:

close, ish just 160x240

QueenMidnight16:

38400?

MidniteSunny:

mhm

QueenMidnight16:

is that all

MidniteSunny:

well ig chu could put it in a sentence at the end, "The area of the room is 38400 ft^2." hope this helped and is right !!!! blessed b

QueenMidnight16:

thank you so much!

QueenMidnight16:

and thanks for tolerating me and my dumb mind that shuts down at the end of the day XD